OpenStudy (anonymous):

Find the maximum value of f(x) = x³ - 6ax² + 9a²x - 2a³ in the interval 00. I will draw my work, pls help

OpenStudy (anonymous):

Actually, the interval is \[0 \le x \le 2\]

OpenStudy (anonymous):

I did this: f'(x) = 3x² - 12ax + 9a² => 3(x² - 4ax + 3a²) => 3(x-3a)(x-a)

OpenStudy (kinggeorge):

Looks good so far.

OpenStudy (anonymous):

K, I also did a table.

OpenStudy (kinggeorge):

From there, recall that local maximums and minimums are at the zero's of the derivative. So the next step would be to set your derivative equal to 0.

OpenStudy (anonymous):

Check this

OpenStudy (anonymous):

I don't know if it's understandable lol

OpenStudy (kinggeorge):

What are the possible solutions for x to \[3(x-3a)(x-a)=0\;?\]

OpenStudy (anonymous):

Those I draw. 0, a and 3a.

OpenStudy (kinggeorge):

Ah, I see what you had written down now. First, note that 0 can't be a solution unless \(a\) was also 0, but \(a>0\) so 0 isn't a solution. So using your solutions of \[x=a\]\[x=3a\]plug those solutions back into the original equation for x.

OpenStudy (anonymous):

In f'(x) or f(x)?

OpenStudy (kinggeorge):

f(x), since we're looking for maximum of f(x).

OpenStudy (anonymous):

I already draw that. When x =a, f(x) = 2a³. When x = 3a, f(x) =0. Check the table.

OpenStudy (anonymous):

I also did it with f'(x) to see the negatives and positives

OpenStudy (kinggeorge):

Well, since \(a>0\) we know that \(2a^3>0\) so \(x=2a^3\) is a maximum.

OpenStudy (anonymous):

I have to know the maximum when a>2

OpenStudy (anonymous):

I guess I need to draw the graph, pellet :/

OpenStudy (kinggeorge):

why when \(a>2\)? Isn't it in the interval \(0\leq x\leq2\)?

OpenStudy (phi):

you should find y'' so you can decide if x=a or x=3a are max or mins.

OpenStudy (anonymous):

Yea, but to find the values in that interval I have to first find when a>2 and all those others. I can't explain it without the graph

OpenStudy (kinggeorge):

What phi said is actually a great way to do it. Find the second derivative, and then plug in the values of x where the first derivative was 0. If it's negative, then you have a maximum, and if it's positive you have a minimum.

OpenStudy (anonymous):

I know that, but the problem is the interval

OpenStudy (anonymous):

The example says I should do the graph

OpenStudy (kinggeorge):

I think I see the issue. You also need to check what happens when \(x=0, 2\) because those might be the max and min in the interval.

OpenStudy (anonymous):

Yea

OpenStudy (kinggeorge):

If you let x=0, you get \(f(0)=-2a^3\) and since \(a>0\), \(a^3>0\) so \(f(0)\) can't be a maximum.

OpenStudy (anonymous):

You're not understanding what I mean, actually. Check my table. I already did that, now I need to find the values in those intervals

OpenStudy (anonymous):

I hate this exercise lol

OpenStudy (kinggeorge):

Well, If you try it with \(x=2\), you find \(f(2)=8-24a+18a^2-2a^3\).

OpenStudy (kinggeorge):

Quick question, is \(a>0\) or \(a>2\)?

OpenStudy (anonymous):

a>0. But in the example a was also > 0 and they solved using a>2, a<2<4a and 4a<2, and the found values were like f(2), f(a), etc

OpenStudy (phi):

break the problem into 0 < a ≤ 2 where f(a) will be a maximum 2<a where f(2) will be the maximum

OpenStudy (anonymous):

I tried to take a pic of the example but my cam sucks lol

OpenStudy (anonymous):

How do I do that, phi?

OpenStudy (anonymous):

Check those graphs and the alternate places where I have to put x so I can find the values in the interval

OpenStudy (anonymous):

King, you there?

OpenStudy (anonymous):

Notice that the first graph is for a>2

OpenStudy (phi):

You already have (almost) break the problem into 0 < a ≤ 2 plug in a into f(x) that is the answer 2<a where f(2) will be the maximum

OpenStudy (kinggeorge):

In the chart, the top row is x, second row is f'(x), and third row is f(x) correct?

OpenStudy (phi):

King already evaluated f(2) up above

OpenStudy (anonymous):

Yes, King

OpenStudy (anonymous):

I can't see that, phi

OpenStudy (anonymous):

I guess I already did that

OpenStudy (anonymous):

lol i'm confused now

OpenStudy (phi):

Think of it this way. If a is less than 2, where is the maximum of the function going to be

OpenStudy (phi):

You know f(a) is a maximum.

OpenStudy (phi):

On the other hand, if a>2 the function is rising toward f(a) so f(2) (the boundary of your interval) must the max

OpenStudy (anonymous):

Oh, that's right!

OpenStudy (anonymous):

But what about the others?

OpenStudy (phi):

The question is Find the maximum value of f(x) = x³ - 6ax² + 9a²x - 2a³ in the interval 0<x<2. Assume a>0. You did: 0<a≤2 2a^3 2<a -2a^3+18a^2-24a+8

OpenStudy (anonymous):

But you're confusing the intervals. When you draw the graph and you alternate positions of a, you have to find the values in those intervals as well

OpenStudy (anonymous):

I will give the example. Let f(x) be x³ - 6ax² + 9a²

OpenStudy (anonymous):

for the same interval

OpenStudy (anonymous):

and a>0 as well

OpenStudy (anonymous):

It would go like: when a>2, max is f(2) = 18a² -24a + 8 when a \[\le\] 2 \[\le\] 4a, that is, 1/2 \[\le\] a \[\le\] 2 (notice both intervals) then max is f(a) = 4a³ when 4a<2, that is, 0<a<1/2, max is f(2) again

OpenStudy (anonymous):

I meant \[a \le 2 \le 4a\] and \[1/2 \le a \le 2\]

OpenStudy (phi):

in this case f(3a) is a minimum, f(a) is a max. a>0 (so f(0) can't be the max)

OpenStudy (kinggeorge):

I think finally see what's going on now. And I agree with phi. We have to break the function up into \(0<a\leq2\) and \(2<a\). Then, we can find the maximum of the function in those parameters.

OpenStudy (anonymous):

Problem is I can't see that without graphs

OpenStudy (phi):

OK, I think I see what you are saying